# Harnessing Multimodal Large Language Models for Multimodal Sequential Recommendation

### v2

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated significant potential in the field of Recommendation Systems (RSs). Most existing studies have focused on converting user behavior logs into textual prompts and leveraging techniques such as prompt tuning to enable LLMs for recommendation tasks. Meanwhile, research interest has recently grown in multimodal recommendation systems that integrate data from images, text, and other sources using modality fusion techniques. This introduces new challenges to the existing LLM-based recommendation paradigm which relies solely on text modality information. Moreover, although Multimodal Large Language Models (MLLMs) capable of processing multi-modal inputs have emerged, how to equip MLLMs with multi-modal recommendation capabilities remains largely unexplored. To this end, in this paper, we propose the Multimodal Large Language Model-enhanced Multimodal Sequential Recommendation (MLLM-MSR) model. To capture the dynamic user preference, we design a two-stage user preference summarization method. Specifically, we first utilize an MLLM-based item-summarizer to extract image feature given an item and convert the image into text. Then, we employ a recurrent user preference summarization generation paradigm to capture the dynamic changes in user preferences based on an LLM-based user-summarizer. Finally, to enable the MLLM for multi-modal recommendation task, we propose to fine-tune a MLLM-based recommender using Supervised Fine-Tuning (SFT) techniques. Extensive evaluations across various datasets validate the effectiveness of MLLM-MSR, showcasing its superior ability to capture and adapt to the evolving dynamics of user preferences.
@@ Introduction
The development of Large Language Models (LLMs) has significantly enhanced the capacity for natural language understanding (Floridi and Chiriatti 2020; Achiam et al. 2023; Touvron et al. 2023), which has been instrumental in advancing recommendation systems (RSs). LLMs have demonstrated remarkable improvements in processing complex user preferences due to its strong semantic understanding and summarization ability. These attributes significantly enhance personalization and accuracy in recommendations(Wu et al. 2023; Zhang et al. 2024a; Ren et al. 2024), particularly in Sequential Recommendations (SRs) where extracting long historical preferences is crucial (Hou et al. 2023; Zheng et al. 2024; Zhai et al. 2023; Li et al. 2023).
Simultaneously, beyond solely modeling textual information, there has been a growing interest in leveraging multimodal information (Liu et al. 2023b). Techniques such as multimodal fusion and gated multimodal units have been utilized to integrate data from various sources—images, videos, and audio—enriching the context for recommendations. This offers a deeper understanding of user-item interactions and naturally leads to the exploration of Multimodal Large Language Models (MLLMs) for enhancing multimodal recommendation systems (Zhang and Tong 2020; Liu et al. 2021b; Zhou and Miao 2024; Kim et al. 2024). MLLMs merge multimodal information into a unified textual semantic space, enhancing the system’s ability to understand and interpret complex data inputs, thereby can significantly improving recommendation accuracy (Liu et al. 2024c; Zhang et al. 2024b). The application of MLLMs in sequential recommender systems presents a promising avenue for dynamically adapting to user preferences and handling the intricate interplay of multimodal data, which holds considerable untapped potential.
@@
Sequential Recommenders (SRs) have progressed from matrix-based models to sophisticated neural architectures. Initially, Factorizing Personalized Markov Chains (FPMC) integrated matrix factorization with Markov chains to model sequential behavior (Rendle, Freudenthaler, and Schmidt-Thieme 2010). The transition to neural models started with GRU4Rec, which employed gated recurrent units for session-based recommendations (Hidasi et al. 2015). Subsequently, SASRec utilized self-attention mechanisms to handle long-term dependencies in user-item interactions (Kang and McAuley 2018), and BERT4Rec introduced transformers to SRs, significantly enhancing performance with deep bidirectional training (Sun et al. 2019).
@@
The integration of Large Language Models (LLMs) into recommendation systems has been profoundly influenced by foundational models such as BERT (Devlin et al. 2018) and GPT-3 (Brown et al. 2020), which demonstrated the potential of LLMs in processing vast amounts of textual data to understand user behaviors deeply. This foundation has been expanded upon by subsequent models like BERT4Rec (Sun et al. 2019) and innovations such as RLMRec (Ren et al. 2024), which tailor LLM capabilities to generate personalized, context-aware recommendations by analyzing detailed user-item interactions.
In the current landscape, LLM applications in recommendation systems are categorized into three main approaches: embeddings-based, token-based, and direct model applications (Wu et al. 2023; Cao et al. 2024). Embeddings-based applications, such as (Cui et al. 2022; Liu et al. 2024b), use LLMs to extract rich feature representations from item and user data, enhancing the system’s understanding of user preferences. Token-based approaches, highlighted in works like (Zhai et al. 2023), focus on generating tokens that capture semantic meanings and potential user preferences, integrating this data into the recommendation logic. Lastly, direct model applications (Hou et al. 2024; Geng et al. 2022) involve using LLMs as end-to-end solutions where the models directly generate recommendations based on user queries and profiles, offering a streamlined and potentially more powerful system architecture. Additionally, there has been an emergence of multimodal LLM-based recommendation frameworks designed to handle scenarios involving multimodal information. These frameworks integrate and process diverse data types such as images, text, and video to enhance the recommendation system’s accuracy and user experience (Liu et al. 2024c; Zhang et al. 2024a).
@@
Our prompt design adheres to foundational methods of multimodal information fusion. Item information can be separated into textual descriptions and image. Hence, In the initial phase, distinct prompts (i.e., text summarization and image description prompt) are used to guide Multimodal Large Language Models (MLLMs) to process these modalities independently, to ensure a more thorough comprehension and detailed feature extraction from each modality, ensuring nuanced characteristics often missed in unified analyses are captured. To ensure both modalities contribute equally to item modeling, the outputs of text summarization and image description are calibrated to similar lengths.
After independently analyzing each modality, our design integrates insights from both textual and visual information using a fusion prompt. This approach aligns with traditional multimodal recommendation strategies that emphasize synthesizing diverse data types to create a comprehensive item profile, enhances the multifaceted understanding of the item (Baltrušaitis, Ahuja, and Morency 2018; Huang, Xiao, and Yu 2019; Zhang and Tong 2020).
@@
After gathering user preferences using the methods described above, we can propose a supervised fine-tuning of an open-sourced Multimodal Large Language Model, such as LLaVA [^1] https://huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf, which excels at understanding images. This model would be used to build a multimodal recommender system through SFT. In line with the definition of sequential recommendation, given a user-item interaction, the MLLM-based recommender system utilizes the prompt contained the obtained user preferences, the textual description and image of the given item and the designed system instruction prompt to predict the probability that the user will interact with the candidate item. Specifically, the prompt designed for the tuned MLLM recommender module is illustrated in Figure 3, where we restrict the output to only include ’yes’ or ’no’ to avoid irrelevant information about the predicted label. Thus, the probability of item interaction can be calculated from the probability score of the predicted first new token as follows:
@@
To construct a multimodal sequential recommender system based on Multimodal Large Language Models (MLLMs), we implement supervised fine-tuning to optimize the model parameters. This fine-tuning process involves adjusting the model to minimize the discrepancy between predicted and actual user interactions. Our dataset construction strategy employs negative sampling, a commonly used training technique in recommendation systems, wherein each positive user-item interaction is coupled with multiple negative samples representing items with which the user did not interact. This methodology aids the model in distinguishing between relevant and irrelevant items through contrastive learning, thereby improving its predictive accuracy.
@@
To assess the performance of baseline methods and our proposed MLLM-MSR for multimodal sequential recommendations, we employed AUC, HR@5, and NDCG@5 as evaluation metrics. To ensure a fair comparison, we standardized the size of the candidate item sets across all baseline methods and our approach.

### v5

@@ Abstract
Recent advances in Large Language Models (LLMs) have demonstrated significant potential in the field of Recommendation Systems (RSs). Most existing studies have focused on converting user behavior logs into textual prompts and leveraging techniques such as prompt tuning to enable LLMs for recommendation tasks. Meanwhile, research interest has recently grown in multimodal recommendation systems that integrate data from images, text, and other sources using modality fusion techniques. This introduces new challenges to the existing LLM-based recommendation paradigm which relies solely on text modality information. Moreover, although Multimodal Large Language Models (MLLMs) capable of processing multi-modal inputs have emerged, how to equip MLLMs with multi-modal recommendation capabilities remains largely unexplored. To this end, in this paper, we propose the Multimodal Large Language Model-enhanced Multimodal Sequential Recommendation (MLLM-MSR) model. To capture the dynamic user preference, we design a two-stage user preference summarization method. Specifically, we first utilize an MLLM-based item-summarizer to extract image feature given an item and convert the image into text. Then, we employ a recurrent user preference summarization generation paradigm to capture the dynamic changes in user preferences based on an LLM-based user-summarizer. Finally, to enable the MLLM for multi-modal recommendation task, we propose to fine-tune a MLLM-based recommender using Supervised Fine-Tuning (SFT) techniques. Extensive evaluations across various datasets validate the effectiveness of MLLM-MSR, showcasing its superior ability to capture and adapt to the evolving dynamics of user preferences. The code of our work is publicly available at https://github.com/YuyangYe/MLLM-MSR.
@@ Introduction
The development of Large Language Models (LLMs) has significantly enhanced the capacity for natural language understanding (Floridi and Chiriatti 2020; Achiam et al. 2023; Touvron et al. 2023), which has been instrumental in advancing recommendation systems (RSs). LLMs have demonstrated remarkable improvements in processing complex user preferences due to its strong semantic understanding and summarization ability. These attributes significantly enhance personalization and accuracy in recommendations(Wu et al. 2023; Zhang et al. 2024; Ren et al. 2024), particularly in Sequential Recommendations (SRs) where extracting long historical preferences is crucial (Hou et al. 2023; Zheng et al. 2024; Zhai et al. 2023; Li et al. 2023).
Simultaneously, beyond solely modeling textual information, there has been a growing interest in leveraging multimodal information (Liu et al. 2023b). Techniques such as multimodal fusion and gated multimodal units have been utilized to integrate data from various sources—images, videos, and audio—enriching the context for recommendations. This offers a deeper understanding of user-item interactions and naturally leads to the exploration of Multimodal Large Language Models (MLLMs) for enhancing multimodal recommendation systems (Liu et al. 2021b; Zhou and Miao 2024; Kim et al. 2024). MLLMs merge multimodal information into a unified textual semantic space, enhancing the system’s ability to understand and interpret complex data inputs, thereby can significantly improving recommendation accuracy (Liu et al. 2024c; Zhang et al. 2024). The application of MLLMs in sequential recommender systems presents a promising avenue for dynamically adapting to user preferences and handling the intricate interplay of multimodal data, which holds considerable untapped potential.
@@
Sequential Recommenders (SRs) have progressed from matrix-based models to sophisticated neural architectures. Initially, Factorizing Personalized Markov Chains (FPMC) integrated matrix factorization with Markov chains to model sequential behavior (Rendle et al. 2010; Yu et al. 2023). The transition to neural models started with GRU4Rec, which employed gated recurrent units for session-based recommendations (Hidasi et al. 2015). Subsequently, SASRec utilized self-attention mechanisms to handle long-term dependencies in user-item interactions (Kang and McAuley 2018), and BERT4Rec introduced transformers to SRs, significantly enhancing performance with deep bidirectional training (Sun et al. 2019).
@@
The integration of Large Language Models into recommendation systems has been profoundly influenced by foundational models such as BERT (Devlin et al. 2018) and GPT-3 (Brown et al. 2020), which demonstrated the potential of LLMs in processing vast amounts of textual data to understand user behaviors deeply. This foundation has been expanded upon by subsequent models like BERT4Rec (Sun et al. 2019) and innovations such as RLMRec (Ren et al. 2024), which tailor LLM capabilities to generate personalized, context-aware recommendations by analyzing detailed user-item interactions.
In the current landscape, LLM applications in recommendation systems are categorized into three main approaches: embeddings-based, token-based, and direct model applications (Wu et al. 2023; Cao et al. 2024; Guo et al. 2024). Embeddings-based applications, such as (Cui et al. 2022; Liu et al. 2024b), use LLMs to extract rich feature representations from item and user data, enhancing the system’s understanding of user preferences. Token-based approaches, highlighted in works like (Zhai et al. 2023), focus on generating tokens that capture semantic meanings and potential user preferences, integrating this data into the recommendation logic. Lastly, direct model applications (Hou et al. 2024; Geng et al. 2022) involve using LLMs as end-to-end solutions where the models directly generate recommendations based on user queries and profiles, offering a streamlined and potentially more powerful system architecture. Additionally, MLLM-based recommendation frameworks have emerged to handle scenarios involving diverse data types like images, text, and video, improving system accuracy and user experience (Liu et al. 2024c; Zhang et al. 2024).
@@
Our prompt design adheres to foundational methods of multimodal information fusion. Item information can be separated into textual descriptions and image. Hence, In the initial phase, distinct prompts (i.e., text summarization and image description prompt) are used to guide MLLMs to process these modalities independently, to ensure a more thorough comprehension and detailed feature extraction from each modality, ensuring nuanced characteristics often missed in unified analyses are captured. To ensure both modalities contribute equally to item modeling, the outputs of text summarization and image description are calibrated to similar lengths.
After independently analyzing each modality, our design integrates insights from both textual and visual information using a fusion prompt. This approach aligns with traditional multimodal recommendation strategies that emphasize synthesizing diverse data types to create a comprehensive item profile, enhances the multifaceted understanding of the item (Baltrušaitis, Ahuja, and Morency 2018; Huang, Xiao, and Yu 2019; Gao et al. 2020).
@@
After gathering user preferences using the methods described above, we can propose a supervised fine-tuning of an open-sourced MLLM, such as LLaVA [^1] https://huggingface.co/llava-hf/llava-v1.6-mistral-7b-hf, which excels at understanding images. This model would be used to build a multimodal recommender system through SFT. In line with the definition of sequential recommendation, given a user-item interaction, the MLLM-based recommender system utilizes the prompt contained the obtained user preferences, the textual description and image of the given item and the designed system instruction prompt to predict the probability that the user will interact with the candidate item. Specifically, the prompt designed for the tuned MLLM recommender module is illustrated in Figure 3, where we restrict the output to only include ’yes’ or ’no’ to avoid irrelevant information about the predicted label. Thus, the probability of item interaction can be calculated from the probability score of the predicted first new token as follows:
@@
To construct a multimodal sequential recommender system based on MLLMs, we implement supervised fine-tuning to optimize the model parameters. This fine-tuning process involves adjusting the model to minimize the discrepancy between predicted and actual user interactions. Our dataset construction strategy employs negative sampling, a commonly used training technique in recommendation systems, wherein each positive user-item interaction is coupled with multiple negative samples representing items with which the user did not interact (Yu et al. 2018). This methodology aids the model in distinguishing between relevant and irrelevant items through contrastive learning, thereby improving its predictive accuracy.
@@
To assess the performance of baseline methods and our proposed MLLM-MSR for multimodal sequential recommendations, we employed AUC, HR@5, and MRR@5 as evaluation metrics (Yu et al. 2020). To ensure a fair comparison, we standardized the size of the candidate item sets across all baseline methods and our approach.